OpenStudy (anonymous):

log(-x)+log(3)=log(2x-5)

OpenStudy (anonymous):

log(-x)+log(3)=log(-3x)

OpenStudy (anonymous):

Log(a) + Log(b) = Log(a*b) Use this: Log(-3x) = Log(2x - 5) Can you solve further??

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

log(2x-5)=log(-3x) so 2x-5 = -3x you surely can solve this ;)

OpenStudy (anonymous):

kk